1A: After​ graduation, you face a choice. One option is to work for a multinational consulting firm and earn a starting salary​(benefits included) of $40,000. The other option is to use $6,000 in savings to start your own consulting firm. You could earn an interest return of 5 percent on your savings. You choose to start your own consulting firm. At the end of the first ​year, you add up all of your expenses and revenues. Your total includes $10,000 in rent, $1,000 in office supplies, $22,000 for office staff, and $5,000 in telephone expenses.

Your explicit costs are $____.

Fill in the blank.

1B: An individual leaves a college​ faculty, where she was earning $70,000 a year, to begin a new venture. She invests her savings of $42,000, which were earning 7 percent annually. She then spends $18,000 renting office equipment, hires two students at $24,000 a year each, rents office space for $9,000, and has other variable expenses of $40,000. At the end of the year, her revenues are $185,000.

Her accounting profit is $____.

Fill in the blank.
Answer #1

1A. Explicit costs = 10,000 + 1,000 + 22,000 + 5,000 = $38,000

       Implicit costs = 40,000 + (6,000 * 0.05) = $40,300

1B. Here, Explicit costs = 18,000 + (2 * 24,000) + 9,000 + 40,000 = $115,000

                Implicit costs = 70,000 + (42,000 * 0.07) = $72,940

Her accounting profit = Total revenue - Explicit costs = 185,000 - 115,000 = $70,000

Her Economic profit = Total revenue - Explicit costs - Implicit costs = 185,000 - 115,000 - 72,940 = -$2,940 (it is a loss)